Calculate the molar solubility of Mg(OH)2 in 1MNH4Cl

KspMg(OH)2 = 1.8 \times 10–11

Kb(NH3) = 1.8 \times 10–5

 

Grade:11

1 Answers

Arun
6 years ago
Dear Student
 
Mg(OH)2(s) +2NH4+  ===== Mg++  + 2NH4OH
1M                     x                       2x
 
let molar solubility of Mg(OH)2 be x in 1 M NH4Cl
 
1.8 * 10^-11 / 1.8*1.8*10^-10    = x* (2x)^2 / 1^2 = 4 x^3/1 = 0.1/1.8 = 1/18
x = 0.24 M
 
regards
Arun (askIITians forum expert)